OpenStudy (anonymous):

The x-intercept of a line is -5 and the y-intercept of the line is -2. What is the equation of the line? o y = - 5/2x - 5 o y = 2/5x + 2 o y = 5/2x -2 o y = -2/5x - 2 How would I find what the b is? Since the formula is y = mx + b.

OpenStudy (anonymous):

Use....Intercept form.. \[\frac{ x }{ a }+\frac{ y }{ b } = 1\]

OpenStudy (anonymous):

a= x inetrcept b = y intercept

OpenStudy (anonymous):

I think it would be the third choice; y = 5/2x - 2. Am I correct?